OpenStudy (anonymous):

Identify the solution(s) of the system of equations, if any. 4x + 2y = 9 y = 5 - 3x

OpenStudy (anonymous):

So I assume you have: 1. 4x + 2y = 9 2. y = 5 - 3x These are simultaneous equations so you can substitute or use elimination so solve them. In this equation they already give us y = ... so we can substitute equation 2 into equation 1. Therefore, we get 4x + 2(5-3x) = 9 4x + 10 - 6x = 9 -2x + 10 = 9 -2x = -1 x = 1/2 Now you can substitute x = 1/2 into equation 1 or 2 to find y. I'll substitute into 2. y = 5 - 3(1/2) = 7/2

OpenStudy (anonymous):

Check the attachment

OpenStudy (anonymous):

Thanks so much everyone! I am just having trouble with re-writing the equation in order to solve for either x or y.